ABCD is a square and P,Q are the midpoint of BC,CD respectively. If AP=a and AQ=b,find in terms of a and b,the directed line segments of AB,AD,BD and AC


1
Expert's answer
2021-02-02T04:32:32-0500

We have that

"a=AB+BC/2.......(1)\\
b=AD+DC/2.......(2)"

But, AD=BC and AB=DC

So,

"a=DC+BC/2\\
b=BC+DC/2"

Solving the simultaneous equation, we have that:

"DC=\frac{2}{3}(2a-b), BC=\frac{2}{3}(2b-a)"

Also,

"AB=DC=\frac{2}{3}(2a-b)\\
AD=BC=\frac{2}{3}(2b-a)"

"AC=AB+BC=\frac{2}{3}(a+b)\\
BD=AD-AB=2(b-a)"